OpenStudy (anonymous):

The average of the test scores of a group of x students is 76 and the average of the test scores of a group of y students is 90. When the scores of the two groups of the students are combined, the average test score is 85. What is the value of x/y?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

5/9

OpenStudy (anonymous):

76x+90y / (x+y) = 85 so u have 5y=9x or x/y=5/9

OpenStudy (kropot72):

The total scores of the x students is 76x and the total scores of the y students is 90y. The combined total scores is 85(x + y). This leads to the following equation: 85x + 85y = 76x + 90y 9x = 5y

OpenStudy (kropot72):

Now divide both sides by 9. Then divide both sides of the result by y.

OpenStudy (anonymous):

@kropot72 Thank you :)

OpenStudy (kropot72):

You're welcome :)